OpenStudy (anonymous):

Find the solution to the following equation by transforming it into a perfect square trinomial. 3x2 + 24x = 27

OpenStudy (anonymous):

3x2 + 24x - 27 = 0 3(x2 + 8x - 9) = 0 x2 + 8x - 9 = 0 (x2 + 8x +____) - 9 = 0 To figure out what goes in the ____ take 1/2 * 8 and square it. That's 4^2 or 16 (x2 + 8x +16) - 9 -16 = 0 (x + 4)2 - 25 = 0 (x + 4)2 = 25 x + 4 = 5 or x + 4 = -5 x = -1 or x = -9

OpenStudy (anonymous):

Thanks babe